OpenStudy (anonymous):

If 5-2x is f(x) what is f'(x)

OpenStudy (anonymous):

f'(x) is pretty much the slope of the equation. so for 5-2x the slope is -2. f'(x)=-2. say if f(x)=5-2x+3(x^2), then f'(x)=6x-2 and to find the slope at any given point plug x in to the formula. but for yours slope is constant throughout from x=(-infinity,infinity)

OpenStudy (anonymous):

f'(x) means 1st derivative of the function f(x), which would be -2.

OpenStudy (anonymous):

the derivative of f(x) is -2, since that is the slope of f(x). The definition of f'(x) is the slope of f(x).

OpenStudy (anonymous):

f'(x) is the slope of f(x)? this is through if x,f(x) represented in Cartesian coordinate sytem. if x & f represent polar coordinates, for instance, \[x = \theta & f = r = r(\theta)\], then \[r'(\theta)\] is no longer be a slope of \[r(\theta)\]. To sum up, we have to know what do variables mean graphically before we decide whether the first derivative of the function is exactly the slope or something else!.

OpenStudy (anonymous):

do you know how to diffrentiate ? or just draw the curve of f 9x) and find the slope. i think you can do it..

OpenStudy (anonymous):

f'(x) or d/dx just gives you the slope. also what the guy above said.. XD

OpenStudy (anonymous):

2 possible answers. Centre of chord joining the 2 points is the origin. The centre of the circle lies on thr perpendicular bisector of this chord. The chord has gradeint 1 and therefore the perpendicular has gradient -1 and passes through the origin. Therefore the centre of the circle lies on the line y=-x and has coordinates of the form(a,-a). The distance of this point from either of the original points must be 4. Using Pythagoras this gives an equation in a which has 2 solutions (a=2 and a=-2). This leads to the 2 answers using ''(x-a)^2 + (y-b)^2=r^2''.